OpenStudy (anonymous):

How do the graphs of y =1/x and y = 5/x+6 compare?

OpenStudy (anonymous):

y=1/x has a vertical asymptote at 0.

OpenStudy (amorfide):

y=1/x draw it and you get an asymptote at y=0 if you draw y=5/x+6 well first of all it is stretched by a factor of 5 so it is steeper then it is translated -6 units in the x axis, meaning it passes through the point -5 not 1